OpenStudy (anonymous):

I need Chemistry help (Connections Unit 6 Lesson 3) 1. A sample of gold (Au) has a mass of 40.14 g. a. Calculate the number of moles of gold (Au) in the sample and record in Table 1. Show your work. b. Calculate the number of atoms of gold (Au) in the sample and record in Table 1. Show your work. 2. A sample of table sugar (sucrose, C12H22O11) has a mass of 1.604 g. a. Calculate the number of moles of C12H22O11 contained in the sample and record in Table 1. Show your work.

OpenStudy (jfraser):

what you need is the relationship between the \(mass\) of a substance and the \(amount\) of that substance, or its quantity. Thankfully, the Periodic Table is full of these relationships. Do you know how to find them?

OpenStudy (awkwardpanda):

I need help with this too. Let me know if you got it.

OpenStudy (anonymous):

number of mole =wight/ molecular wight =(40.14)/196.96=0.2 mole

OpenStudy (anonymous):

number of atoms= 6.02214129×10^23*0.2=1.2*10^22

OpenStudy (anonymous):

C12H22O11 molecular weight of C = 12*12=144 H= 22*1=22 O= 11*16=176 molecular weight of sugar= sum of all =144+22+176=342 number of mole = 1.604/342=0.00469

OpenStudy (anonymous):

*weight number of mole =weight/ molecular weight =(40.14)/196.96=0.2 mole

OpenStudy (awkwardpanda):

For A. I got: 40.14gx1mole/196.97g=0.2037 and for b i got 0.2037 * 6.022x10^23= 1.2266x10^23. is that right?

OpenStudy (jfraser):

that's it